Mapa mental
PROBABILITY
Contenido del mapa
PROBABILITY
VALUE
- Sum of Probabilities of all elementary events is 1. For events A,B,C; P(A)+P(B)+P(C)= 1.
THEORETICAL PROBABILITY
- 0<P(E) <1 P(E)= Number of trials in which the event happened/Total no. of trials.
EXPERIMENTAL PROBABILITY
- What actually happens in an experiment: P(E) = Number of outcomes favourable to E/ No. of all possible outcomes of the experiment.
EXAMPLES
- Dice : Two dice are rolled, what is the probability of getting 12 as a sum? Solution:No. of possible outcomes=36 No. of favourable outcomes=1 i.e. (6,6) P(E)=1/36
- Card: What is the probability of getting an ace from a pack of 52 cards? Solution: No. of favourable outcomes = 4 No. of possible outcomes = 52, P(E) = 4/52 = 1/13
- Coin: When a coin is tossed, what would be the probability of appearing head? Solution: Total outcomes = 2 i.e.(H,T) Favourable outcomes = 1 Probability P(E) = 1/2
DEFINITIONS
- Complementary Event: For event E, complement event, P(E) = 1-P(E)
- Elementary Event: Event having only one outcome of the experiment.
- Sure or Certain Event: When an event having probability to occur as 1.
Información
- Descripción
- Mapa Mental sobre PROBABILITY, creado por Aishika Bhowmik el 08/11/2021.
- Creado por
- Aishika Bhowmik
- Visitas
- 0
- Valoración